OpenStudy (anonymous):

How many solutions are there to the following system of equations? 4x – 14y = 6 –2x + 7y = –3 A. infinitely many B. 2 C. 0 D. 1

OpenStudy (dtan5457):

4x – 14y = 6 –2x + 7y = –3 multiply that 2nd equation by 2 to get -4x+14y=-6 add that to 4x – 14y = 6 what do you notice about the answer?

OpenStudy (anonymous):

is it A.?

OpenStudy (dtan5457):

Yes, they are the same line, therefore they have infinite solutions

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

What is the solution to the system of equations? 5x+5y=8 x-2y=10

OpenStudy (anonymous):

4/5 ooor 0.8

OpenStudy (anonymous):

guys?